November 2000: Status report

On Nov. 1, the so-called "1200 m cavity" has been locked for the first time. This cavity consists of the power-recycling mirror and one of the interferometer arms. This step took longer than expected, since the windows of the vacuum gate valves caused spatial beam distortions although they had been ordered with very high optical specifications. After pumping down the central cluster and opening the gate valves, the distortions disappeared and the cavity could be locked.Next steps: Over the following two months we will test this cavity and optimize the laser frequency stabilization, measure alignment noise and implement the automatic alignment of the power recycling cavity. Further the electrostatic actuator for the differential armlength control of the Michelson interometer will be tested.

Main Suspensions
The triple pendulum suspensions for the second Michelson arm (the "north arm") will be finished parallel to this, and the installation procedure for the first monolithic suspension will be tested. The pulling machine for the silica fibers has been transferred from Glasgow to Hannover. Geppo has trained some people in Hannover to produce perfect fibers. We expect to install the first monolithic-suspended end mirror in January in the north end building.

Data Acquisition
The data acquisition system is in place and was debugged. It will be used by the scientist on the site to test the installed subsystems. The last missing part in the DAQ chain will be uploaded and tested at the end of November. The part in question is the frame writer that converts the internal GEO file format into the common "FRAME" format used by the other GW projects.

Laser System
The laser is subject to a long-term test in the University lab. No power degradation could be seen for over 60 days. The system will be installed at the site in late December.

December 2000: Status report

The installation of the mode-cleaner system is a success. The mode-cleaner lock works automatically overnight together with the auto-alignment. It runs very regularly with only one reset within fourteen hours. Now we are studying the long-term drift in order to obtain data for a better optimization. The whole system has been stabilized on the power-recycling cavity.

The suspension for the beamsplitter is ready. All the installations in the north-end building are finished, up to the control program (network problems).

Merry Xmas & a Happy New Year to All of the Gravity Wave Community! 

 
loading content
Go to Editor View